OpenStudy (anonymous):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (texaschic101):

3/5 cups of apple juice + 3/4 cups of cranberry juice = 3/5 + 3/4 = 12/20 + 15/20 = 27/20 total juice set up a proportion.. 3/(27/20) = 1/x --- (3 servings to 27/20 juice = 1 serving to x juice) cross multiply (3)(x) = (27/20)(1) 3x = 27/20 x = (27/20) / 3 x = 27/20 * 1/3 x = 27/60 when reduced = 9/20 check... 9/20 * 3 = 27/20 27/20 = 27/20 (correct) There are 9/20 cups of juice per serving :)